A cup of coffee (100 g) cools from 90˚C down to comfortable room temperature 25.˚C. How much energy does it release to the surroundings?


1
Expert's answer
2021-09-28T07:38:25-0400

The energy release during cooling is


"Q=cm(t_f-t_i)=4200·0.1(25-90)=27300\text{ J}."